OpenStudy (anonymous):

What is the vertex for the graph of y = 3x^2 − 12x + 9?

OpenStudy (anonymous):

(2,-3)

OpenStudy (anonymous):

the 3rd eqn is a quad.

OpenStudy (anonymous):

x = 3/2 and x = -1/2

OpenStudy (anonymous):

x=-5 and x=-6

OpenStudy (anonymous):

What are the approximate solutions of 4x2 + 3 = −12x to the nearest hundredth?